128 - 188 x (57 - 282) - 585 = ?
1
Expert's answer
2012-08-23T07:15:30-0400
128 - 188 x (57 - 282) - 585=128 - 188 x ( - 225) - 585=128+188*225-585=-457+188*225=41843